# Asymptotic Safety and Canonical Quantum Gravity

### v1

@@ 3 Towards candidate observables for quantum gravity
With CQG being formulated in the Hamiltonian framework, the evolution of spacetime is governed by constraints. These are the Hamiltonian constraint and the momentum constraint, which generate time evolution and spatial diffeomorphisms (see Ref. 76 for a detailed analysis of Hamiltonian constrained systems). Following the quantum mechanical prescription, a physical observable is then an operator which commutes with all constraints. This is the definition of a Dirac observable [7, 8, 9, 77]. Such physical observables must be invariant under the gauge symmetries generated by the constraints, including both spatial and temporal coordinate transformations.
@@
Importantly, in AS one must distinguish between the RG scale, which sets the coarse-graining resolution in the RG flow, and the physical momentum appearing in scattering processes or in the propagators. While the RG flow governs how effective couplings evolve under a change of the RG scale, this does not directly furnish the running wrt. the momentum of external states. It is the running with respect of the external momenta which represents the physical running [96, 97, 98]. How to connect these two runnings is a current active topic of investigation. A comparison between these two different runnings and their qualitative agreement in the graviton propagator can be found in Refs. 99, 100.
@@
On the other hand, while CQG comes with a well-defined notion of observables, there is no fully established systematic mechanism for transitioning to an effective description in the IR regime. Moreover, its notion of “effective” dynamics in general differs from the Lagrangian field theoretical one. Among the attempts to introduce coarse-graining in CQG-related approaches we mention coarse-graining in spin foams via tensor network renormalization [117, 118, 119, 120, 121, 122] , also used to study perfect discretizations [123, 124, 125], and in Tensorial Group Field Theory (TGFT) [126, 127]. Furthermore, Hamiltonian renormalization methods have been recently advanced, focusing on constructing the correct continuum (ultraviolet) Hamiltonian via RG flows [128, 129, 130]. It is important to emphasize that these advancements are aimed at the construction of the continuum limit. Remarkably, effective spin foams have been recently developed, which allow computational access to the continuum limit [131, 132] and only recently the obtained continuum limit has been analyzed via FRG methods [133]. Finally, there has been important progress in constructing the graviton propagator from spin foams [134, 135, 136, 137, 138, 139] and in accessing the low energy perturbative regime by means of a large-spin expansion [140, 141, 142].
@@
In particular, within the procedure sketched in (3), several works have shown that the non-perturbative quantum dynamics of background-independent approaches (especially in Loop Quantum Cosmology (LQC) [163, 164] and Group Field Theory (GFT) [165, 166]) can be recast as classical modified-gravity actions. In exceptional cases one inverts the effective equations from the quantum Hamiltonian (Legendre transform) to find a covariant action. In others a suitable modified Lagrangian is postulated whose field equations reproduce the same dynamics as the effective Hamiltonian. As an example, a unique Palatini $f(R)$ action whose field equations exactly reproduce the isotropic LQC bounce has been derived [167]. Starting from the LQC effective Friedmann equation for a massless scalar, one can solve for a connection-independent Lagrangian $f(R)$ that yields the same dynamics. The resulting covariant action (an infinite series in $R$) leads to exactly the non-singular bounce of LQC. Subsequently, it was shown that metric $f(R)$ and higher-order gravity theories can reproduce LQC dynamics, treating higher-derivative terms as perturbative corrections around GR [168]. Later works extended this method to more general curvature invariants and modified LQC models, deriving families of higher curvature actions that mimic the LQC bounce phenomenologically [169]. Independently, mimetic and degenerate higher-order scalar–tensor theories, particularly the Chamseddine–Mukhanov’s limiting-curvature model [170], were shown to exactly reproduce LQC’s bounce equations in isotropic settings, offering an alternative and more precise correspondence [159, 154].
Always within (3), among LQG covariant path integral quantum gravity models, spin foam models offer a path-integral formulation of quantum gravity [24, 171, 172], from which one can extract an effective, semiclassical Lagrangian. It has been shown that if the vertex amplitude has the right large-spin behavior, the path integral reproduces Regge calculus in the semiclassical limit [173, 174, 175].
@@ 3.2 Gravitational Hamiltonian and gravitational Lagrangian
It is a longstanding open issue, how the Lagrangian and the Hamiltonian formalism can be related in a gravitational context (see appendix E.2 in Ref. 184). This problem is indeed rooted in the need of defining a time in the Hamiltonian context.
@@
where $N$ is the lapse function, $N^{i}$ the shift vector, and $q_{ij}$ the $(d-1)$ spatial metric on constant time slices. The action becomes
@@
where ${}^{(d-1)}R[q]$ is the $(d-1)$-dimensional curvature constructed from $q_{ij}$ and $K_{ij}$ the extrinsic curvature
@@
While the relational formalism introduces dynamics by selecting a preferred temporal reference, the Lagrangian approach preserves full covariance and instead handles the redundancy associated with gauge symmetries through a gauge-fixing procedure and the introduction of ghosts à la Faddeev–Popov [185, 186] is needed, in order to account for the unphysical degrees of freedom. Gauge-fixing independence has been proven to hold in a perturbative loop expansion, imposing order by order on-shell conditions [53]. In this setting, gauge-invariant observables are obtained from evaluating correlation functions or amplitudes on-shell. However their construction typically involves the choice of a background. Conversely, in non-perturbative settings such as in AS, the results are still affected by gauge-parameter dependence [187, 103, 104, 105]. Furthermore, the regularization method chosen also generally affects the full quantum gauge invariance [26]. In order to alleviate this problem, within AS the form factor program has been established [188, 189, 190, 191, 192]: the idea is to study the non-local effective action for gravity (and matter fields) using form factors, which are Laplacian- or momentum-dependent functions appearing in front of curvature invariants (or other operators). Form factors encode the scale-dependent (and generally non-local) structure of quantum corrections. This representation allows one to keep background gauge invariance manifest throughout the RG flow, which contrasts with ansätze based on local operators in the Einstein–Hilbert or polynomial expansions that often introduce artificial gauge or parametrization dependence. Because form factors resemble quantities that appear in on-shell amplitudes, they bring the framework closer to settings where gauge invariance is manifest.
@@
A closely related issue which has been source of concern is the choice of the measure. While typically from a Lagrangian point of view, the measure is not uniquely fixed, provided it satisfies the given symmetry conditions, from a Hamiltonian (relational) perspective, the construction of the physical configuration space and its conjugate momentum space depends on the choice of clock and is hence fixed. This might give rise to differences in local measure factors, discussed since the 1960s in geometrodynamics and background-dependent quantizations [198, 199, 200, 201]. These were argued to contribute only to divergent terms in higher-loop amplitudes and to be dependent on the chosen regularization scheme. While certain regularization choices can cancel or suppress these contributions, the ambiguity surrounding their role has been recently been clarified [202]. These factors are however essential, as they properly reflect the off-shell symmetries generated by the constraints [193, 203]. Importantly, when relating the Hamiltonian framework to the Lagrangian path integral, the measure derived from the conjugate variables is also fixed and it may vary depending on the chosen physical frame and might be highly non-trivial. We will come back to this in Section 6.
@@ 4.1 The Functional Renormalization Group
[…]
@@ 4.2 RG flows of gravity
where $h$ is the fluctuation as specified in equation (1). Evaluating the trace in the flow equation then boils down to compute the trace over the index structure and over the eigenvalues of the d’Alembert operator ¯$\overline{\square}\equiv\bar{g}_{\mu\nu}\bar{\nabla}^{\mu}\bar{\nabla}^{\nu}$.
@@
[…]
@@
$(\mathcal{Q}_{\text{ FP}})^{\mu}\,_{\nu}c^{\nu}\equiv-\sqrt{16\pi G_{N}(k)}\mathcal{L}_{c}F^{\mu}=-\left(\bar{g}^{\mu\lambda}\bar{g}^{\nu\rho}-\frac{1}{2}\bar{g}^{\nu\mu}\bar{g}^{\rho\lambda}\right)\bar{\nabla}_{\nu}(g_{\rho\sigma}\nabla_{\lambda}c^{\sigma}+g_{\lambda\sigma}\nabla_{\rho}c^{\sigma})\,,$ (37)
@@
Having a positive $\text{Re}\;\theta_{1,2}>0$, the two operators in the Einstein–Hilbert truncation are associated to two relevant directions, i.e., related perturbations grow under the RG flow toward the IR. Furthermore, the critical exponents being complex conjugate is a manifestation of the spiraling of the trajectories around the UV fixed point [258]. For more details we refer the reader to Refs. 259, 28.
@@
At this stage let us mention that there are several excellent reviews and books on asymptotic safety, including Refs. 25, 26, 260, where more details and applications are reported. The program has found more evidence for such a fixed point, by refining the ansatz for $\Gamma_{k}$ [261, 262], including field redefinitions [263] within the essential RG [264], by coupling different matter fields to gravity [265, 266], by using different carrier fields than just the metric fluctuation [267, 268] and by studying bi-metric flows [56]. Important to mention is that asymptotic safety can also be tested via alternative RG methods [255, 194] or in a discrete setting by lattice gravity, as it is done in dynamical triangulations [58, 269] (for connections between discrete and continuous realizations we refer to Ref. 270, 271, 236, 272) . As a side development aside from non-perturbative methods, evidence has been furnished, about the realization of the non-trivial RG UV fixed point within perturbation theory, by using background independent techniques, such as the proper time flow or dimensional regularization in a non-minimal subtraction scheme [273, 194].
@@
The issue of the theory’s signature has long been a cause for concern. It is only in the last couple of years that flourishing attempts to perform a Lorentzian computation have been pushed forward and are now well-established [274, 275, 276, 100, 277, 278, 279, 105, 280, 281]. The issue of going Lorentzian is twofold: on one side there is the technical challenge to compute the trace as described above for the Lorentzian flow equation; on the other side, several interpretational issues regarding the meaning of the cutoff arise due to the ambiguity of regularizing spacelike and timelike momenta (no well-defined integrating-out ordering is given). As for the first challenge, among the most promising methods, it has been proposed to exploit methods well established in the Algebraic QFT framework [278] on one side, and to extend the heat kernel (or better said Schrödinger) traces to Lorentzian signature on the other side (see Refs. 282, 248, 283 for a connection between the two methods and [284, 285] for a discussion on the Schrödinger group from an analytically continued lapse). In fact, recently, a way how to solve the resulting flow equation by means of a cutoff in heat kernel space has been pushed forward [73, 281, 286]. The heat kernel expansion in this context reads:
@@
Hence, we notice that the price to pay for such a Lorentzian heat kernel expansion is that the coupling constants become generically complex. For a discussion about this new approach we refer the reader to Refs. 73, 281. What the second Lorentzian issue is concerned, such an ambiguity is unavoidable but it also sheds light on some interpretational aspects which in Euclidean signature were hidden. For instance, it seems natural to think that the selection of which modes have to be integrated out is dependent on the experimental setting. This would introduce an observer dependence into the RG process, invalidating any possibility of having a unique Lorentzian flow. This was investigated in the context of de Sitter spacetime by analyzing the running spectral flow [65] and opens the avenue for more interesting investigations on physical, relational aspects of the RG methodology in Lorentzian signature.

### v2

@@ 2 Shared principles, different strategies
In addition to this, it should be noted that in this review both AS and CQG are formulated within a quantization paradigm, assuming a pre-existing classical spacetime structure to be quantized. In contrast, a number of other approaches take an emergent perspective, positing that the gravitational field is a collective phenomenon arising from more fundamental, possibly non-spatiotemporal, degrees of freedom. Examples include Group Field Theory (GFT) [74, 75, 76], causal set theory [77, 78], and also, in a certain sense, canonical LQG and lattice-based formulations [79]. In all these frameworks, the fundamental variables are not smooth fields but discrete or combinatorial structures, from which the continuum spacetime of GR can arise only through collective dynamics [80].
A related conceptual commonality concerns spacetime topology change. While recent continuum and discrete studies [81, 82, 83, 84, 85, 86, 87] have investigated quantum gravitational dynamics allowing for topology change, both CQG and the AS-RG formulation usually assume fixed (globally hyperbolic) topologies.
@@ 3 Towards candidate observables for quantum gravity
With CQG being formulated in the Hamiltonian framework, the evolution of spacetime is governed by constraints. These are the Hamiltonian constraint and the momentum constraint, which generate time evolution and spatial diffeomorphisms (see Ref. 90 for a detailed analysis of Hamiltonian constrained systems). Following the quantum mechanical prescription, a physical observable is then an operator which commutes with all constraints. This is the definition of a Dirac observable [7, 8, 9, 91]. Such physical observables must be invariant under the gauge symmetries generated by the constraints, including both spatial and temporal coordinate transformations.
@@
Importantly, in AS one must distinguish between the RG scale, which sets the coarse-graining resolution in the RG flow, and the physical momentum appearing in scattering processes or in the propagators. While the RG flow governs how effective couplings evolve under a change of the RG scale, this does not directly furnish the running wrt. the momentum of external states. It is the running with respect of the external momenta which represents the physical running [110, 111, 112]. How to connect these two runnings is a current active topic of investigation. A comparison between these two different runnings and their qualitative agreement in the graviton propagator can be found in Refs. 113, 114. Furthermore, recent work has suggested that the distinction between standard and physical running may depend on the choice of renormalization point, and that any running of the gravitational couplings might only manifest in curved spacetime, where curvature sets a physical scale [115].
@@
On the other hand, while CQG comes with a well-defined notion of observables, there is no fully established systematic mechanism for transitioning to an effective description in the IR regime. Moreover, its notion of “effective” dynamics in general differs from the Lagrangian field theoretical one. Among the attempts to introduce coarse-graining in CQG-related approaches we mention coarse-graining in spin foams via tensor network renormalization [132, 133, 134, 135, 136, 137] , also used to study perfect discretizations [138, 139, 140], and in Tensorial Group Field Theory (TGFT) [141, 142]. In particular, within TGFT, both perturbative and FRG methods have been developed to explore the continuum and non-perturbative limits. The perturbative approach has been formulated directly in the language of spin-foam amplitudes, showing that these correspond to a specific truncation or sector of the complete TGFT dynamics [143, 144]. The FRG formulation, on the other hand, provides a more general framework, allowing for the identification of fixed points and critical behavior [145, 146]. This makes possible, at least in principle, a systematic comparison of continuum-limit properties, such as critical exponents. In addition, mean-field and condensate methods have been applied to TGFTs, providing a technically simplified way to probe continuum, collective regimes [147, 148, 149].
Furthermore, Hamiltonian renormalization methods have been recently advanced, focusing on constructing the correct continuum (ultraviolet) Hamiltonian via RG flows [150, 151, 152]. It is important to emphasize that these advancements are aimed at the construction of the continuum limit. Remarkably, effective spin foams have been recently developed, which allow computational access to the continuum limit [153, 154] and only recently the obtained continuum limit has been analyzed via FRG methods [155].
Finally, a number of studies have explored the reconstruction of the graviton propagator from spin foam amplitudes [156, 157, 158, 159, 160, 161] and the access to a low-energy, perturbative regime through large-spin expansions [162, 163, 164]. While these investigations have provided valuable insights into how linearized gravitational correlations could, in principle, emerge from spin foam dynamics, it must be emphasized that they rely on extremely simple simplicial complexes and very limited degrees of freedom. As such, they probe only a small portion of the full theory’s dynamics and remain far from testing the continuum limit, which is essential for any meaningful notion of graviton propagation. In this respect, a substantial open challenge for the spin foam and canonical frameworks alike lies in demonstrating how continuum perturbative gravity can emerge from the nonperturbative discrete dynamics.
@@
In particular, within the procedure sketched in (3), several works have shown that the non-perturbative quantum dynamics of background-independent approaches (especially in Loop Quantum Cosmology (LQC) [185, 186] and GFT [74, 76]) can be recast as classical modified-gravity actions. In exceptional cases one inverts the effective equations from the quantum Hamiltonian (Legendre transform) to find a covariant action. In others a suitable modified Lagrangian is postulated whose field equations reproduce the same dynamics as the effective Hamiltonian. As an example, a unique Palatini $f(R)$ action whose field equations exactly reproduce the isotropic LQC bounce has been derived [187]. Starting from the LQC effective Friedmann equation for a massless scalar, one can solve for a connection-independent Lagrangian $f(R)$ that yields the same dynamics. The resulting covariant action (an infinite series in $R$) leads to exactly the non-singular bounce of LQC. Subsequently, it was shown that metric $f(R)$ and higher-order gravity theories can reproduce LQC dynamics, treating higher-derivative terms as perturbative corrections around GR [188]. Later works extended this method to more general curvature invariants and modified LQC models, deriving families of higher curvature actions that mimic the LQC bounce phenomenologically [189]. Independently, mimetic and degenerate higher-order scalar–tensor theories, particularly the Chamseddine–Mukhanov’s limiting-curvature model [190], were shown to exactly reproduce LQC’s bounce equations in isotropic settings, offering an alternative and more precise correspondence [181, 176].
Always within (3), among LQG covariant path-integral quantum gravity models, spin foam models offer a path integral formulation of quantum gravity [24, 191, 192], from which one can extract an effective, semiclassical Lagrangian. It has been shown that, if the vertex amplitude exhibits the correct large-spin behavior, the spin-foam path integral reproduces Regge calculus in this discrete semiclassical limit [193, 194, 195]. While this represents an important consistency check, it should be emphasized that reproducing classical Regge calculus on a fixed simplicial complex is not equivalent to recovering continuum GR. The correspondence concerns the discrete classical regime, not the continuum limit, and several open issues remain even at the level of the large-$j$ approximation [196, 137, 140, 197]. Indeed, as illustrated by the experience with simplicial quantum gravity approaches such as quantum Regge calculus and causal or dynamical triangulations [198, 70], establishing a genuine continuum limit from a discrete quantum theory poses substantial conceptual and technical challenges, despite the fact that discrete models trivially reproduce classical Regge calculus by construction.
@@
This comparison also highlights a deeper conceptual distinction between the two frameworks. In LQG, coarse-graining does not simply correspond to a cutoff in modes of the same kind, as in the AS RG-based approach, but rather to a transition between qualitatively different kinds of microscopic structures. Even if the discrete geometries underlying LQG were regarded as regularization devices, the reconstruction of a smooth spacetime would still amount to an emergent, collective phenomenon.
Related to this point, it is important to distinguish clearly between semiclassical and continuum approximations or limits. The former concern the recovery of classical dynamics from quantum operators on suitable semiclassical states, while the latter refer to the emergence of smooth spacetime behavior from the collective dynamics of fundamentally discrete or combinatorial structures. These two notions are conceptually and technically distinct, and the continuum limit, rather than the semiclassical one, is what captures the genuine coarse-grained, collective dynamics discussed above. Clarifying this distinction is essential when comparing the effective descriptions arising in approaches such as LQG, spin foam models, or GFT with those obtained through the functional RG in AS.
@@ 3.2 Gravitational Hamiltonian and gravitational Lagrangian
It is a longstanding open issue, how the Lagrangian and the Hamiltonian formalism can be related in a gravitational context (see appendix E.2 in Ref. 207). This problem is indeed rooted in the need of defining a time in the Hamiltonian context.
@@
where $N$ is the lapse function, $N^{i}$ the shift vector, and $q_{ij}$ the (d1)(d-1) spatial metric on constant time slices. The action becomes
@@
where $^{(d-1)}R[q]$ is the (d1)(d-1)-dimensional curvature constructed from $q_{ij}$ and $K_{ij}$ the extrinsic curvature
@@
While the relational formalism introduces dynamics by selecting a preferred temporal reference, the Lagrangian approach preserves full covariance and instead handles the redundancy associated with gauge symmetries through a gauge-fixing procedure and the introduction of ghosts à la Faddeev–Popov [208, 209] is needed, in order to account for the unphysical degrees of freedom. In this context, gauge-fixing independence has been proven to hold only in a perturbative loop expansion, imposing order by order on-shell conditions [53]. In this setting, gauge-invariant observables are obtained from evaluating correlation functions or amplitudes on-shell. However their construction typically involves the choice of a background. Conversely, in non-perturbative settings such as in AS, the results are still affected by gauge-parameter dependence [210, 118, 119, 120]. Furthermore, the regularization method chosen also generally affects the full quantum gauge invariance [26]. In order to alleviate this problem, within AS the form factor program has been established [211, 212, 213, 214, 215]: the idea is to study the non-local effective action for gravity (and matter fields) using form factors, which are Laplacian- or momentum-dependent functions appearing in front of curvature invariants (or other operators). Form factors encode the scale-dependent (and generally non-local) structure of quantum corrections. This representation allows one to keep background gauge invariance manifest throughout the RG flow, which contrasts with ansätze based on local operators in the Einstein–Hilbert or polynomial expansions that often introduce artificial gauge or parametrization dependence. Because form factors resemble quantities that appear in on-shell amplitudes, they bring the framework closer to settings where gauge invariance is manifest.
@@
It should be noted at this stage, that the introduction of a “preferred” reference in the relational strategy refers to a physical frame, typically implemented through dynamical reference fields, and not to a coordinate choice or gauge-fixing. When formulated in this way, the relational construction can be implemented entirely at the Lagrangian level while maintaining full diffeomorphism invariance of the resulting observables, as exemplified in generally covariant relational frameworks (non-reduced phase space) and in applications such as TGFT cosmology [147, 148, 76]. This was also the path followed in the construction of Ref. 117. In principle, therefore, the relational formalism is perfectly compatible with general covariance: the introduction of reference fields merely provides a way to parametrize physical observables in a diffeomorphism-invariant theory. What changes is the representation of dynamics, expressed relationally with respect to the chosen physical frame, rather than the symmetry content of the theory. In practice, however, concrete implementations often involve approximations that single out a particular physical frame or neglect the backreaction and quantum properties of the reference system. In such cases, covariance may appear broken or only approximately realized, even though the underlying formalism remains covariant. The challenge is thus not a conceptual incompatibility with general covariance, but the technical problem of embedding relationally defined, on-shell observables into an off-shell covariant quantization framework.
@@
A closely related issue which has been source of concern is the choice of the measure. While typically from a Lagrangian point of view, the measure is not uniquely fixed, provided it satisfies the given symmetry conditions, from a Hamiltonian (relational) perspective, the construction of the physical configuration space and its conjugate momentum space depends on the choice of clock and is hence fixed. This might give rise to differences in local measure factors, discussed since the 1960s in geometrodynamics and background-dependent quantizations [221, 222, 223, 224, 225]. These were argued to contribute only to divergent terms in higher-loop amplitudes and to be dependent on the chosen regularization scheme. While certain regularization choices can cancel or suppress these contributions, the ambiguity surrounding their role has been recently been clarified [226]. These factors are however essential, as they properly reflect the off-shell symmetries generated by the constraints [216, 227]. Importantly, when relating the Hamiltonian framework to the Lagrangian path integral, the measure derived from the conjugate variables is also fixed and it may vary depending on the chosen physical frame and might be highly non-trivial. We will come back to this in Section 6.
@@ 3.3 Remarks on more empirical connections between gravitational Hamiltonian and Lagrangian
It is also important to note that current spin foam models and their GFT formulations differ structurally from canonical LQG in a crucial way. The covariant models are rooted in simplicial geometry, employing piecewise-flat building blocks already in the definition of the quantum amplitudes [238, 229, 233, 239, 24]. This simplicial setting restricts the underlying graphs or complexes to fixed valence and makes it difficult to establish a clear correspondence with the canonical quantum dynamics, in particular with a Hamiltonian constraint formulated for smooth geometries [240, 241]. Although various generalizations at the combinatorial level have been proposed to bring spin foam and GFT frameworks closer to the canonical theory [241, 242, 243], their amplitudes remain tied to the simplicial construction and lose their geometric justification when extended beyond it. This mismatch is reflected mathematically in the failure of cylindrical consistency, a key feature of the kinematical Hilbert space in canonical LQG, which is not satisfied by any of the current spin foam or GFT models [244, 137, 140, 197].
@@ 4.1 The Functional Renormalization Group
$G_{k}(p)\sim\frac{1}{Z_{k}p^{2}}\overset{\text{at fixed point}}{\sim}\frac{1}{p^{2-\eta}}\;.$ (26)
@@ 4.2 RG flows of gravity
where $h$ is the fluctuation as specified in equation (1). Evaluating the trace in the flow equation then boils down to compute the trace over the index structure and over the eigenvalues of the d’Alembert operator $\overline{\square}\equiv\bar{g}_{\mu\nu}\bar{\nabla}^{\mu}\bar{\nabla}^{\nu}$.
@@
$+\frac{1}{2}\left(E+\frac{1}{6}\bar{R}\;\mathbbm{1}\right)^{2}+\frac{1}{12}[\bar{\nabla}_{\mu},\nabla_{\nu}][\bar{\nabla}^{\mu},\bar{\nabla}^{\nu}]+\frac{1}{6}\bar{\nabla}^{2}\;\left(E+\frac{1}{6}\bar{R}\;\mathbbm{1}\right)\;.$
@@
[…]
@@
Having a positive $\text{Re}\;\theta_{1,2}>0$, the two operators in the Einstein–Hilbert truncation are associated to two relevant directions, i.e., related perturbations grow under the RG flow toward the IR. Furthermore, the critical exponents being complex conjugate is a manifestation of the spiraling of the trajectories around the UV fixed point [289]. For more details we refer the reader to Refs. 290, 28.
@@
At this stage let us mention that there are several excellent reviews and books on asymptotic safety, including Refs. 25, 26, 291, where more details and applications are reported. The program has found more evidence for such a fixed point, by refining the ansatz for $\Gamma_{k}$ [292, 293], including field redefinitions [294] within the essential RG [295], by coupling different matter fields to gravity [296, 297], by using different carrier fields than just the metric fluctuation [298, 299] and by studying bi-metric flows [56]. Important to mention is that asymptotic safety can also be tested via alternative RG methods [286, 217] or in a discrete setting by lattice gravity, as it is done in dynamical triangulations [58, 79] (for connections between discrete and continuous realizations we refer to Ref. 300, 301, 267, 302) . As a side development aside from non-perturbative methods, evidence has been furnished, about the realization of the non-trivial RG UV fixed point within perturbation theory, by using background independent techniques, such as the proper time flow or dimensional regularization in a non-minimal subtraction scheme [303, 217].
@@
The issue of the theory’s signature has long been a cause for concern. It is only in the last couple of years that flourishing attempts to perform a Lorentzian computation have been pushed forward and are now well-established [304, 305, 306, 114, 307, 308, 309, 120, 310, 311]. The issue of going Lorentzian is twofold: on one side there is the technical challenge to compute the trace as described above for the Lorentzian flow equation; on the other side, several interpretational issues regarding the meaning of the cutoff arise due to the ambiguity of regularizing spacelike and timelike momenta (no well-defined integrating-out ordering is given). As for the first challenge, among the most promising methods, it has been proposed to exploit methods well established in the Algebraic QFT framework [308] on one side, and to extend the heat kernel (or better said Schrödinger) traces to Lorentzian signature on the other side (see Refs. 312, 279, 313 for a connection between the two methods and [314, 315] for a discussion on the Schrödinger group from an analytically continued lapse). In fact, recently, a way how to solve the resulting flow equation by means of a cutoff in heat kernel space has been pushed forward [73, 311, 316]. The heat kernel expansion in this context reads:
@@
Hence, we notice that the price to pay for such a Lorentzian heat kernel expansion is that the coupling constants become generically complex. For a discussion about this new approach we refer the reader to Refs. 73, 311. What the second Lorentzian issue is concerned, such an ambiguity is unavoidable but it also sheds light on some interpretational aspects which in Euclidean signature were hidden. For instance, it seems natural to think that the selection of which modes have to be integrated out is dependent on the experimental setting. This would introduce an observer dependence into the RG process, invalidating any possibility of having a unique Lorentzian flow. This was investigated in the context of de Sitter spacetime by analyzing the running spectral flow [65] and opens the avenue for more interesting investigations on physical, relational aspects of the RG methodology in Lorentzian signature.